40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11925     Entry time: Mon Jan 11 19:01:56 2016     In reply to: 11924     Reply to this: 11926
Author: gautam 
Type: Update 
Category: LSC 
Subject: PLL Marconi Investigation 

EDIT 01/12/2016 6PM: I've updated the plots of the in-loop spectra such that they are calibrated throughout the entire domain now. I did so by inferring the closed-loop transfer function (G/(1-G)) from the measured open-loop transfer function (G), and then fitting the inferred TF using vectfit4 (2 poles). The spectra were calibrated by multiplying the measured spectra by the magnitude of the fitted analytic TF at the frequency of interest.

EricQ brought back one of the Marconis that was borrowed by the Cryo lab to the 40m today (it is a 2023B - the Marconi used for all previous measurements in this thread was 2023A). Koji had suggested investigating the frequency noise injected into the PLL by the Marconi, and I spent some time investigating this today. We tried to mimic the measurement setup used for the earlier measurements as closely as possible. One Marconi was used as a signal source, the other as the LO for the PLL loop. All measurements were done with the carrier on the signal Marconi set to 310MHz (since all our previous measurements were done around this value). We synced the two Marconis by means of the "Frequency Standard" BNC connector on the rear panel (having selected the appropriate In/Out configurations digitally first). Two combinations were investigated - with either Marconi as LO and signal source. For each combination, I adjusted the FM gain on the Marconi (D in the plot legends) and the overall control gain on the SR560 (G in the plot legends) such that their product remained approximately constant. I measured the PLL OLG at each pair to make sure the loop shape was the same throughout all trials. Here are the descriptions of the attached plots:

Attachment #1: 2023A as LO, 2023B as source, measured OLGs

Measured OLG for the various combinations of FM gain and SR560 gain tested. The UGF is approximately 30kHz for all combinations - the exceptions being D 1.6MHz, G=1e4 and D=3.2MHz, G=1e4. I took the latter two measurements just because these end up being the limiting values of D for different carrier frequencies on the Marconi.

Attachment #2: 2023A as LO, 2023B as source, measured spectra of control signal (uncalibrated above 30kHz)

I took the spectra down to 2Hz, in two ranges, and these are the stitched versions. 

Attachment #3: 2023B as LO, 2023A as source, measured OLGs

Attachment #4: 2023B as LO, 2023A as source, measured spectra of control signal (uncalibrated above 30kHz)

So it appears that there is some difference between the two Marconis? Also, if the frequency noise ASD-frequency product is 10^4 for a healthy NPRO, these plots suggest that we should perhaps operate at a lower value of D than the 3.2MHz/V we have been using thus far? 

As a quick trial, I also took quick spectra of the PLL control signals for the PSL+Aux X and PSL+Aux Y beat signals, with the 2023B as the LO (Attachment #5). The other difference is that I have plotted the spectrum down to 1 Hz (they are uncalibrated above 30Hz). The PSL+Y combination actually looks like what I would expect for an NPRO (for example, see page 2 of the datasheet of the Innolight Mephisto) particularly at lower frequencies - not sure what to make of the PSL+X combination. Also, I noticed that the amplitude of the PSL+Y beatnote was going through some large-amplitude (beat-note fluctuates between -8dBm and -20dBm) but low frequency (period ~10mins) oscillations. This has been observed before, not sure why its happening though. 

More investigations to be done later tonight.

Attachment 1: 2023ALockedto2023B.pdf  74 kB  Uploaded Tue Jan 12 18:07:51 2016  | Hide | Hide all | Show all
2023ALockedto2023B.pdf
Attachment 2: 2023ALockedto2023B_spectra.pdf  98 kB  Uploaded Tue Jan 12 18:07:58 2016  | Hide | Hide all | Show all
2023ALockedto2023B_spectra.pdf
Attachment 3: 2023BLockedto2023A.pdf  73 kB  Uploaded Tue Jan 12 18:08:06 2016  | Show | Hide all | Show all
Attachment 4: 2023BLockedto2023A_spectra.pdf  99 kB  Uploaded Tue Jan 12 18:08:13 2016  | Hide | Hide all | Show all
2023BLockedto2023A_spectra.pdf
Attachment 5: TestSpectra.pdf  68 kB  Uploaded Tue Jan 12 18:08:21 2016  | Hide | Hide all | Show all
TestSpectra.pdf
Attachment 6: 2016_01_AUXLaser.tar.gz  673 kB  Uploaded Tue Jan 12 18:08:31 2016
ELOG V3.1.3-